OpenStudy (anonymous):

Write an equation of the line that passes through the given points. (-8,5) and (1,0) The equation is

OpenStudy (dan815):

equation of a line is

OpenStudy (dan815):

y=mx+b or y-y1=m(x-x1)

OpenStudy (dan815):

m=slope= change in y/change in x

OpenStudy (anonymous):

y = - (5/9)x -1